Discrete Probability Distribution

Probability Mass Function

Denoted by PP. Gives the probability that a discrete random variable XX is exactly equal to some value xx.

∑xP(x)=1\sum_{x} P(x) = 1

Parameters

Mean

E(X)=μ=∑ixi⋅P(X=xi)E(X) = \mu = \sum_{i} x_i \cdot P(X = x_i)
Here:
xix_i
each possible value of XX
P(X=xi)P(X = x_i)
the probability of observing that value

Variance

Var(X)=σ2=∑i(xi−μ)2⋅P(X=xi)\text{Var}(X) = \sigma^2 = \sum_{i} (x_i - \mu)^2 \cdot P(X = x_i)

An equivalent computational formula is:

Var(X)=∑ixi2⋅P(X=xi)−μ2\text{Var}(X) = \sum_{i} x_i^2 \cdot P(X = x_i) - \mu^2

Cumulative Distribution Function

F(x)=∑t≤xP(X=t)F(x) = \sum_{t \leq x} P(X = t) F(x)=F(⌊x⌋)    for x∈ZcF(x) = F(\lfloor x \rfloor)\;\;\text{for } x \in \mathbb{Z}^{c}

Example

Consider a discrete random variable XX with the following probability distribution:

XP(X)
10.2
20.3
30.4
40.1
Solution Reveal Hide

The mean would be: E(X)=1⋅0.2+2⋅0.3+3⋅0.4+4⋅0.1=2.4E(X) = 1 \cdot 0.2 + 2 \cdot 0.3 + 3 \cdot 0.4 + 4 \cdot 0.1 = 2.4.

The variance would be: Var(X)=(1−2.4)2⋅0.2+(2−2.4)2⋅0.3+(3−2.4)2⋅0.4+(4−2.4)2⋅0.1=0.84\text{Var}(X) = (1-2.4)^2 \cdot 0.2 + (2-2.4)^2 \cdot 0.3 + (3-2.4)^2 \cdot 0.4 + (4-2.4)^2 \cdot 0.1 = 0.84.

Types

Uniform Distribution

A type of discrete probability distribution where all outcomes are equally likely. If a random variable XX can take nn distinct values, each value has a probability of P(X=x)=1nP(X = x) = \frac{1}{n}.

Bernoulli Distribution

A type of discrete probability distribution where there are only two possible outcomes, often referred to as “success” and “failure”. If a random variable XX can take values 0 or 1, with P(X=1)=pP(X = 1) = p and P(X=0)=1−pP(X = 0) = 1 - p, then XX follows a Bernoulli distribution.

PMF of Bernoulli distribution:

P(x)=px(1−p)1−x    ;  x=0  or  1P(x) = p^x (1-p)^{1-x} \;\;;\; x = 0\;\text{or}\;1
Here:
pp
the probability of success

Binomial Distribution

When an experiment consists of nn independent repeated Bernoulli trials, the experiment is said to be a binomial experiment. The number of successes (XX) follows a binomial distribution. And has a PMF:

P(X)=(nx)px(1−p)n−x    ;  x=0,1,…,nP(X) = \binom{n}{x} p^x (1-p)^{n-x}\;\;;\; x = 0, 1, \ldots, n
Here:
nn
the number of trials
θ\theta
the probability of success

Assumptions made in binomial distribution:

  • Each trial is independent.
  • Each trial has only two possible outcomes.
  • The probability of success is constant across trials.
  • The number of trials is fixed.

For a large sample of a binomial distribution:

μ=np    ∧    σ=np(1−p)\mu = np \;\; \land \;\; \sigma = \sqrt{np(1-p)}

Poisson Distribution

If XX is the number of independent successes occur within a fixed time interval, then XX is said to follow a Poisson distribution. And has a PMF:

P(x)=e−λλxx!    ;  x=0,1,2,…P(x) = \frac{e^{-\lambda} \lambda^x}{x!}\;\;;\; x=0,1,2,\ldots
Here:
λ\lambda

average number of successes occurred within the fixed time interval

E(X)=λ=Var(X)\text{E}(X) = \lambda = \text{Var}(X)

If Xi∼Poisson(λi)X_i \sim \text{Poisson}(\lambda_i) where i=1,2,…,ni = 1, 2, \ldots, n, and XiX_i is independent then:

∑i=1nXi∼Poisson(∑i=1nλi)\sum_{i=1}^n X_i \sim \text{Poisson}\bigg( \sum_{i=1}^n \lambda_i \bigg)
Written by November 3, 2026 2 min read
Was this helpful?